javonarnold:

Ruiz needs to buy a countertop for a kitchen. He calculated the area to be 16 square feet. The actual area is 15.4 square feet. What is Ruiz's percentage error? 0.6% 3.75% 3.9% 6%

Xxbabyjas04xx:

16-15.4=0.6 0.6/15.4=3.89610 3.89610*100%=3.9%

javonarnold:

so 3.9 or0.6

javonarnold:

@KjSaif

KjSaif:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold @KjSaif \(\color{#0cbb34}{\text{End of Quote}}\) 3.9%

javonarnold:

oh ok thanks brodie